OpenStudy (anonymous):

6 more than 1/4 of a number is 1/3 of the number

OpenStudy (anonymous):

Isolate your key words. 6 more implies what operation? 1/4 of a number implys what operation? 1/3 of a number implies what operation? Try to then setup your equation.

OpenStudy (anonymous):

i did set it up i know the answer should be 72 don't know how to show the work.......1/4x+6=1/3x

OpenStudy (anonymous):

not bad but your notation may be a bit off. it should be \[\frac{ x }{ 4 }+6 = \frac{ x }{ 3 }\]

OpenStudy (anonymous):

with the equation setup, you want to get the 'x' terms on one side of the equals sign and the non-x terms on the opposite sign first.

OpenStudy (anonymous):

opposite side* not sign.

OpenStudy (anonymous):

6=x/3-x/4

OpenStudy (anonymous):

good now find a common denominator for the fractions and and combine them.

OpenStudy (anonymous):

6=4x/12-3x/12

OpenStudy (anonymous):

6*12=1x/12*12 72=x

OpenStudy (anonymous):

perfect.

OpenStudy (anonymous):

thanks wanted to make sure i solved it correctly